# Improving Cramér–Rao Bound And Its Variants: An Extrinsic Geometry Perspective

### v1

@@ Abstract
This work presents a geometric refinement of the classical Cramér–Rao bound (CRB) by incorporating curvature-aware corrections based on the second fundamental form associated with the statistical model manifold. That is, our formulation shows that relying on the extrinsic geometry of the square root embedding of the manifold in the ambient Hilbert space comprising square integrable functions with respect to a fixed base measure offers a rigorous (and intuitive) way to improve upon the CRB and some of its variants, such as the Bhattacharyya-type bounds, that use higher-order derivatives of the log-likelihood. Precisely, the improved bounds in the latter case make explicit use of the elegant framework offered by employing the Faà di Bruno formula and exponential Bell polynomials in expressing the jets associated with the square root embedding in terms of the raw scores. The interplay between the geometry of the statistical embedding and the behavior of the estimator variance is quantitatively analyzed in concrete examples, showing that our corrections can meaningfully tighten the lower bound, suggesting further exploration into connections with estimator efficiency in more general situations.
@@ 1 Introduction
The Cramér-Rao Bound (CRB) is a foundational result in parametric estimation theory, specifying a lower limit on the variance of any (locally) unbiased estimator under appropriate regularity conditions, and underpins the concept of estimator efficiency [1, 2, 3]. While widely used, the CRB may be loose in several practical settings – such as nonlinear models, low signal-to-noise regimes, or finite sample sizes – leading to extensive literature on improved lower bounds. Notable classical refinements include the Chapman-Robbins bound [4] and the Ziv-Zakai bound [5], which generalize the CRB beyond the usual regularity assumptions, such as differentiability under the integral sign, or invoke optimal detection problems. Bayesian and hybrid variants of the CRB have been developed to incorporate prior uncertainty that avoid costly matrix inversions [6]. Computational saving was also achieved with constrained CRBs and singular Fisher information matrices in [7].
@@
A line of work that is of relevance to us includes the classical Bhattacharyya inequality ([8]; see also [9, 10]) that generalizes the CRB by incorporating higher-order derivatives of the log-likelihood and is of importance when an estimator that is efficient in the sense of the CRB is not available. While clarifying the algebraic structure of the Bhattacharyya bound, a hierarchy of improvements – an increasing sequence of variance bounds that converge to the estimator variance – was shown by Pillai and Sinha [11] based on projections of the estimator error onto spans of higher-order derivatives of the log-likelihood. This elegant formulation of the Bhattacharyya bound, however, lacks a geometric interpretation despite its Hilbert space foundation. In short, these constructions remain intrinsically score-based and extrinsic geometry of the embedded statistical manifold, which results in nontrivial correction if the estimator error does not lie in the span of the higher-order derivatives considered, is not accounted for; this is exactly what we point out in our research.
Information geometry offers an alternative framework in which statistical models are viewed as Riemannian manifolds endowed with the Fisher-Rao metric. This viewpoint allows estimator performance to be studied using geometric concepts such as curvature, geodesics, and connections [12]. Nevertheless, most information geometry-based works tend to focus on intrinsic properties and divergence measures, as briefly discussed below.
@@
Several works have extended the CRB to intrinsic Riemannian settings. In an important contribution, Smith [13] derived intrinsic versions of the CRB on manifolds utilizing an arbitrary affine connection with arbitrary geodesics for both biased and unbiased estimators while Barrau and Bonnabel [14], in a follow-up, elucidated the extent to which the intrinsic CRB obtained by Smith is independent of a specific choice of coordinates. Further, the authors also addressed a question revolving around the dependence of the intrinsic bound on the underlying parameter. Extending these results to the Bayesian setting, Bouchard et al. [15] studied the problem of covariance matrix estimation when the data is sampled from a Gaussian distribution whose covariance matrix is drawn from an inverse Wishart distribution. Performance bounds for this problem were obtained, for both the mean squared error (Euclidean metric) and the natural Riemannian distance for Hermitian positive definite matrices (affine invariant metric), using the derived intrinsic Bayesian CRB. Interestingly, it was shown that the maximum a posteriori and the minimum mean squared error estimators are asymptotically efficient in the intrinsic case, and not so in the Euclidean situation.
More recently, building upon the work of Li and Zhao [16] who derived the Wasserstein–Cramér-Rao inequality – a lower bound on the Wasserstein variance of an estimator, Nishimori and Matsuda [17] investigated conditions for an estimator to attain the lower bound asymptotically. In an interesting application to chemical reaction networks, Loutchko et al. [18] developed, among various other results, a multivariate CRB through the comparison of two Riemannian metric tensors.
@@
Another important thread involves divergence-based generalizations of the CRB. Mishra and Kumar [19] proposed a general Riemannian metric based on relative $\alpha$-entropy to obtain a generalized Bayesian Cramér-Rao inequality, thereby establishing a lower bound for the variance of an unbiased estimator for the $\alpha$-escort distribution starting from an unbiased estimator for the underlying distribution. Subsequently, Kumar and Mishra [20] provided a unifying differential geometric framework for deriving CRB-like inequalities from generalized Csiszár divergences. They applied Eguchi’s theory to derive the Fisher information metric and dual affine connections arising from these generalized divergence functions, which enabled them to prove a more widely applicable version of the Cramér-Rao inequality. In [21], these authors again relied on Eguchi’s theory of obtaining dualistic geometric structures from a divergence function and then applied Amari-Nagoaka’s theory [12] to obtain a Cramér-Rao type inequality. These approaches serve to find unbiased and efficient estimators, but typically remain in an intrinsic, divergence-induced geometry.
@@
However, we note that a conceptual gap remains. The algebraic projection-based CRBs, such as those of Bhattacharyya [8]/Pillai and Sinha [11], do not exploit the geometric curvature of the statistical manifold. While it may be argued that the second order Bhattacharyya-type bounds also incorporate second derivatives of the density and hence (in some sense) capture second-order information, these do not directly/correctly capture the inherent geometry via the curvature information. Conversely, information geometry-based CRBs focus on intrinsic structures and rarely intersect with projection-theoretic Hilbert space methods. In particular, there is little work incorporating extrinsic geometric effects – such as the second fundamental form – into variance bounds.
@@
In this work, we bridge this gap by developing a Hilbert space projection-based and Riemannian geometric refinement of classical and higher-order CRBs. Suppose the family of probability distributions $\{P_{\theta}\}$, parameterized by the scalar $\theta\in\Theta$, is absolutely continuous with respect to a fixed base measure $\mu$ with a strictly positive density $f(\cdot;\theta)$. Define the square root embedding $s_{\theta}:\Theta\mapsto L^{2}(\mu)$ by $s_{\theta}:=\sqrt{f(\cdot;\theta)}$, where $L^{2}(\mu)$ is the space of square integrable functions with respect to $\mu$, and consider the jets
@@
Starting from a projection-theoretic decomposition of estimator error in the ambient space $L^{2}(\mu)$, we derive a residual component orthogonal to the span of $\eta_{1}$. We then show that the contribution of this residual can be explicitly quantified using the second fundamental form of the statistical manifold embedded in $L^{2}(\mu)$, yielding a curvature-aware correction to the CRB. A variance bound that explicitly betters the Bhattacharyya/Pillai–Sinha-type results is then derived by further analyzing the structure of the $\eta_{k}$ in relation to the true (raw) score and its derivatives $\partial_{\theta}^{k}\log f(\cdot;\theta)$. Importantly, note that we refer to results that use $\partial_{\theta}^{k}\log f(\cdot;\theta)$ for $k\geq 1$ as Bhattacharyya-type (or Pillai–Sinha-type) bounds throughout the paper since the classical Bhattacharyya bound used $\frac{\partial^{k}_{\theta}f}{f}$ instead.
Our contributions are as follows:
  • We derive extrinsic geometric refinement terms using the second fundamental form, providing new insight into how curvature controls estimator error beyond linear approximations, and establish improvement over the classical CRB – see Theorem 2.
  • We interpret the higher-order variance lower bound sequence obtained by Bhattacharyya [8]/Pillai and Sinha [11] as successive $L^{2}$ projections, laying the foundation for geometric corrections making use of higher derivatives of the square root embedding and curvature information “beyond a certain order" (cf. Theorem 5).
  • These basic results, and those of Bhattacharyya/Pillai–Sinha-type, are further strengthened by a careful analysis of the form of the jets $\eta_{k}$ expressed in terms of the raw score and its derivatives $\partial_{\theta}^{k}\log f(\cdot;\theta)$ via the Faà di Bruno formula and the exponential Bell polynomials. Specifically, we make use of a curvature-aware term, akin to that in the CRB case in Theorem 2, along with orthogonal components arising from Faà di Bruno expansions of the embedding; this result can be found in Proposition 8.
  • The proposed framework is exemplified by applying it to explicit models, demonstrating strict (and fully analytical) improvement over both classical CRBs and algebraic higher-order bounds.
To sum up, this work unifies algebraic and geometric approaches to variance lower bounds and represents, to the best of our knowledge, the first use of extrinsic curvature to sharpen the CRB. It opens new avenues in the intersection of information geometry, Hilbert space theory, and statistical estimation.

### v2

@@ Abstract
This work presents a geometric refinement of the classical Cramér–Rao bound (CRB) in the non-asymptotic regime by incorporating curvature-aware corrections based on the second fundamental form associated with the statistical model manifold. That is, our formulation shows that relying on the extrinsic geometry of the square root embedding of the manifold in the ambient Hilbert space comprising square integrable functions with respect to a fixed base measure offers a rigorous (and intuitive) way to improve upon the CRB and some of its variants, such as the Bhattacharyya-type bounds, that use higher-order derivatives of the log-likelihood. Precisely, the improved bounds in the latter case make explicit use of the elegant framework offered by employing the Faà di Bruno formula and exponential Bell polynomials in expressing the jets associated with the square root embedding in terms of the raw scores. The interplay between the geometry of the statistical embedding and the behavior of the estimator variance is quantitatively analyzed in concrete examples, showing that our corrections can meaningfully tighten the lower bound, suggesting further exploration into connections with estimator efficiency in more general situations.
@@ 1 Introduction
The Cramér-Rao Bound (CRB) is a foundational result in parametric estimation theory, specifying a lower limit on the variance of any (locally) unbiased estimator under appropriate regularity conditions, and underpins the concept of estimator efficiency [1, 2, 3]. While widely used, the CRB may be loose in several practical settings – such as nonlinear models, low signal-to-noise regimes, or finite sample sizes – leading to extensive literature on improved lower bounds. Notable classical refinements include the Chapman-Robbins bound [4] and the Ziv-Zakai bound [5], which generalize the CRB beyond the usual regularity assumptions, such as differentiability under the integral sign, or invoke optimal detection problems. Bayesian and hybrid variants of the CRB have been developed to incorporate prior uncertainty that avoid costly matrix inversions [6]. Computational savings were also achieved with constrained CRBs and singular Fisher information matrices in [7].
@@
A line of work that is of relevance to us includes the classical Bhattacharyya inequality ([8]; see also [9, 10]) that generalizes the CRB by incorporating higher-order derivatives of the log-likelihood and is of importance when an estimator that is efficient in the sense of the CRB is not available. Conditions under which a Bhattacharyya bound is a valid lower bound for the variance were studied in [11] by Seth, and considerations regarding attainment of the Bhattacharyya bound can be found in the work of Fend [12]. It is also important for us to note that the Bhattacharyya bound can be viewed as a hierarchy of improvements over the CRB – an increasing sequence of variance bounds that converge to the estimator variance – based on projections of the estimator error onto spans of higher-order derivatives of the (log)likelihood. Convergence of the sequence of bounds to the variance of the best unbiased estimator for certain functions of the underlying parameter was shown in [13] by Blight and Rao when the sampling distribution is one among normal, binomial, Poisson, negative binomial, or gamma with scale parameter. The same result was extended by Ghosh and Sathe for all estimable functions of the parameter and all multiparameter exponential families [14], provided the true value of the parameter is an interior point of the natural parameter space. Similar considerations were also made later by Pillai and Sinha [15]. However, these elegant formulations of the Bhattacharyya bound lack a geometric interpretation despite the Hilbert space foundation. In short, these constructions remain intrinsically score-based and extrinsic geometry of the embedded statistical manifold, which results in nontrivial correction if the estimator error does not lie in the span of the considered higher-order derivatives, is not accounted for; this is exactly what we point out in our research.
Information geometry offers an alternative framework in which statistical models are viewed as Riemannian manifolds endowed with the Fisher-Rao metric. This viewpoint allows estimator performance to be studied using geometric concepts such as curvature, geodesics, and connections [16]. A pioneering work of direct relevance to ours is that of Efron [17] who tried to quantify the proximity of an arbitrary one-parameter family to the standard exponential class of distributions, which is, in some sense, a geodesic in the space of all possible probability distributions. The exponential family enjoys nice statistical properties; for instance, the maximum likelihood estimator (MLE) is a sufficient statistic and, provided that we choose an appropriate function of the parameter to estimate, attains the CRB. Efron introduced the notion of “statistical curvature” of a one-parameter family at a given point and showed that large curvature corresponds to a breakdown of the good properties just alluded to; for a specific example, the variance of the MLE exceeds the CRB in approximate proportion to the squared statistical curvature – asymptotically, the MLE extracts all but a part of the information in a sample that is proportional to the curvature squared. To add some more detail, Efron considered the problem of estimating the scalar parameter in a curved exponential family, on the basis of an i.i.d. sample of size $n$, using a squared error loss function to evaluate consistent and efficient estimators that are smooth functions of the sufficient statistic. There he proved that the $O(\tfrac{1}{n^{2}})$ term in the variance asymptotics precisely involves the squared statistical curvature, linking it to Rao’s influential second-order efficiency theory. It is pertinent to note that Efron’s curvature is the curvature of an embedding associated with the particular way a parameter space is placed inside a natural, higher-dimensional parameter space, as measured by the second fundamental form.
A related fact, in the context of deriving the asymptotic MSE of a bias-corrected first-order efficient estimator, is given in Theorem 4.4 of [16], where the statistical curvature term is identified with the e-curvature of the model manifold. In a similar vein to Efron’s work, Takeuchi and Akahira [18] rigorously analyzed the higher-order efficiency of sequential estimation procedures in the scalar parameter case (their results also hold for nonsequential estimation). Continuing along these lines, Okamoto et al. [19] developed a higher-order asymptotic theory of sequential estimation in the framework of the geometry of multidimensional curved exponential families leading to a design principle of the second-order efficient sequential estimation procedure.
We note with great interest that Jim Reeds, in a discussion on Efron’s work, points out that he concludes statements about the coefficients of the asymptotic expansions of the variance, but not about the variance itself as captured in conclusions using the CRB and Bhattacharyya inequalities. This fact, the square root embedding, and considering the relationship between the log-likelihood and square root embeddings to evaluate the effectiveness, represent three important points of departure in our work from Efron’s significant contribution and the other papers mentioned above in this body of literature.
@@
Most other information geometry-based works tend to focus on intrinsic properties and divergence measures, as briefly discussed below. Several papers have extended the CRB to intrinsic Riemannian settings. In an important contribution, Smith [20] derived intrinsic versions of the CRB on manifolds utilizing an arbitrary affine connection with arbitrary geodesics for both biased and unbiased estimators while Barrau and Bonnabel [21], in a follow-up, elucidated the extent to which the intrinsic CRB obtained by Smith is independent of a specific choice of coordinates. Further, the authors also addressed a question revolving around the dependence of the intrinsic bound on the underlying parameter. Extending these results to the Bayesian setting, Bouchard et al. [22] studied the problem of covariance matrix estimation when the data is sampled from a Gaussian distribution whose covariance matrix is drawn from an inverse Wishart distribution. Performance bounds for this problem were obtained, for both the MSE (Euclidean metric) and the natural Riemannian distance for Hermitian positive definite matrices (affine invariant metric), using the derived intrinsic Bayesian CRB. Interestingly, it was shown that the maximum a posteriori and the minimum MSE estimators are asymptotically efficient in the intrinsic case, but not so in the Euclidean situation. In a slightly different vein, Loutchko et al. [23] developed, among other results, a multivariate CRB by exploring the Riemannian geometrical aspects of the information geometry of chemical reaction networks.
@@
Another important thread involves divergence-based generalizations of the CRB. Mishra and Kumar [24] proposed a Riemannian metric based on relative $\alpha$-entropy to obtain a generalized Bayesian Cramér-Rao inequality, thereby establishing a lower bound for the variance of an unbiased estimator for the $\alpha$-escort distribution starting from an unbiased estimator for the underlying distribution. Subsequently, Kumar and Mishra [25] provided a unifying differential geometric framework for deriving CRB-like inequalities from generalized Csiszár divergences. They applied Eguchi’s theory to derive the Fisher information metric and dual affine connections arising from these generalized divergence functions, which enabled them to prove a more widely applicable version of the Cramér-Rao inequality. In [26], these authors again relied on Eguchi’s theory of obtaining dualistic geometric structures from a divergence function and then applied Amari-Nagoaka’s theory [16] to obtain a Cramér-Rao type inequality. These approaches serve to find unbiased and efficient estimators, but typically remain in an intrinsic, divergence-induced geometry.
@@
However, we note that a conceptual gap remains. The algebraic projection-based CRBs, such as those of Bhattacharyya and some of the follow-ups we met above, do not exploit the geometric curvature of the statistical manifold. While it may be argued that the Bhattacharyya-type bounds following Efron’s paper capture second-order curvature information, these are all asymptotic and do not capture the inherent geometry in non-asymptotic variance expressions. In contrast, many of the information geometry-based CRBs focus on intrinsic structures and rarely intersect with projection-theoretic Hilbert space methods. This motivates our research on incorporating extrinsic geometric effects – such as the second fundamental form – into (non-asymptotic) variance bounds.
In recent years, there has been a rapidly growing body of work that develops estimation theory, efficiency notions, and algorithmic methods based on the Wasserstein geometry of probability measures, complementing the classical Fisher–Rao framework. A central contribution in this direction is due to Li and Zhao [27], who introduced the Wasserstein information matrix (WIM) by pulling back the Otto–Wasserstein metric from the space of probability measures to finite-dimensional parametric statistical models. This construction yields Wasserstein analogues of score functions, covariance operators, and a Wasserstein–CRB, together with asymptotic efficiency results for Wasserstein natural gradient methods. Subsequent work has further clarified the statistical meaning and attainability of these bounds. In particular, Nishimori and Matsuda [28] studied conditions under which estimators attain the Wasserstein–CRB asymptotically, identifying Wasserstein analogues of exponential families and establishing Wasserstein efficiency for location–scale models. Closely related is the work of García Trillos, Jaffe, and Sen [29], who developed a detailed Wasserstein–Cramér–Rao theory for unbiased estimation, interpreting Wasserstein variance as a notion of estimator sensitivity to infinitesimal perturbations of the data, and this insight allowed them to derive a collection of results which are analogous to the classical case; for instance, a characterization of models in which there exist unbiased estimators achieving the lower bound exactly. At a more structural level, Li and Zhao [30] investigate scaling limits of WIM on Gaussian mixture models, showing that suitably renormalized pullback Wasserstein metrics admit well-defined limits as component variances vanish. Their analysis yields generalized Wasserstein metrics on probability simplices, including second-order corrections and associated Wasserstein gradient flows. This line of work deepens the understanding of Wasserstein-induced geometry on parametric statistical models, but is primarily concerned with metric limits and dynamical behavior, rather than variance lower bounds or curvature-induced corrections to classical Fisher-based Cramér–Rao inequalities.
These Wasserstein-based approaches represent a substantial conceptual shift: rather than refining the classical variance bound associated with the Fisher–Rao (or Hellinger) geometry, they redefine the notion of estimator instability using transport-based metrics. As a consequence, the resulting bounds are not directly comparable to the classical Cramér–Rao or Bhattacharyya inequalities, but instead address robustness and sensitivity properties that are invisible to Fisher-based variance analysis.
Complementary to the above Wasserstein–Cramér–Rao–type developments, several works emphasize the role of the ambient metric in shaping representations and induced geometries on families of probability distributions. Arias-Castro and Qiao [31] study the embedding of distributional data via classical scaling and Isomap, adapting tools from Euclidean dimensionality reduction to probability measures and explicitly highlighting how the choice of metric (in particular, Wasserstein-type metrics) governs the resulting geometry and statistical structure. While their focus is on representation and visualization rather than estimation-theoretic bounds, their results further underscore the central role played by the ambient geometry of distribution space.
A related line of work compares Fisher and Wasserstein geometries directly. Amari and Matsuda [32] studied affine deformation models and showed that Wasserstein geometry naturally separates shape and deformation parameters, trading Fisher efficiency for robustness against waveform perturbations; they further demonstrated that Wasserstein estimators coincide with MLE only in the Gaussian case. Fukushi, Nakanishi-Ohno, and Matsuda [33] investigated the intrinsic and extrinsic geometry of location–scale–shape models under the Wasserstein metric, showing that while such models may be intrinsically flat, they can still exhibit extrinsic curvature in the ambient Wasserstein space. These results highlight how extrinsic geometry plays a role in Wasserstein statistics, albeit in a setting fundamentally different from Fisher–Rao variance analysis.
From a computational and algorithmic perspective, Wasserstein geometry has also been extensively studied in connection with optimization, sampling, and gradient flows. Wu and Gardner [34] analyzed stochastic natural gradient variational inference, establishing non-asymptotic convergence rates for conjugate models and clarifying the geometric limitations of natural gradient methods in non-conjugate settings. Zuo’s recent dissertation [35] further advances the computational theory of Wasserstein gradient flows, developing primal–dual damping algorithms, neural approximations of Wasserstein flows, and accelerated Langevin-type samplers. These works emphasize computational efficiency and dynamics in Wasserstein space, rather than statistical lower bounds for estimator variance.
Finally, Lawson et al. [36] provided a comprehensive and practitioner-oriented account of Fisher geometry for multivariate normal distributions, focusing on geodesics, distances, and numerical computation of shortest paths. Their work elucidates the curvature induced by covariance structures and compares Fisher geodesics with alternative paths such as those arising from Wasserstein geometry, but does not address variance lower bounds or projection-based refinements of the Cramér–Rao inequality.
In contrast to the above Wasserstein-based and algorithmic frameworks, the present paper remains entirely within the classical Fisher–Rao variance paradigm and addresses a complementary question: how to sharpen non-asymptotic variance lower bounds for (locally) unbiased estimators by exploiting the extrinsic geometry of the statistical model embedded in an ambient Hilbert space. Our approach does not redefine variance or efficiency, nor does it replace the Fisher information metric. Instead, it refines the classical CRB and Bhattacharyya-type bounds using log-likelihood derivatives by quantifying the contribution of estimator error components orthogonal to the tangent space, using the second fundamental form of the square root embedding/normal components that lie beyond the span of the scores. To the best of our knowledge, this constitutes the first systematic use of extrinsic curvature to obtain explicit, non-asymptotic improvements over the CRB, the emphasis being on the qualification non-asymptotic, thereby bridging projection-theoretic variance bounds and geometric information theory in a manner distinct from both intrinsic Fisher geometry and Wasserstein-based regularizations.
@@
That is, in this work, we bridge the gaps identified above by developing a Hilbert space projection-based and Riemannian geometric refinement of classical and higher-order CRBs. Suppose the family of probability distributions $\{P_{\theta}\}$, parameterized by the scalar $\theta\in\Theta$, is absolutely continuous with respect to a fixed base measure $\mu$ with a strictly positive density $f(\cdot;\theta)$. Define the square root embedding $s_{\theta}:\Theta\mapsto L^{2}(\mu)$ by $s_{\theta}:=\sqrt{f(\cdot;\theta)}$, where $L^{2}(\mu)$ is the space of square integrable functions with respect to $\mu$, and consider the jets
@@
Starting from a projection-theoretic decomposition of the estimator error in the ambient space $L^{2}(\mu)$, we derive a residual component orthogonal to the span of $\eta_{1}$. We then show that the contribution of this residual can be explicitly quantified using the second fundamental form of the statistical manifold embedded in $L^{2}(\mu)$, yielding a curvature-aware correction to the CRB. A higher-order version of this result is also stated. Further, a variance bound that explicitly betters Bhattacharyya-type results using derivatives of the log-likelihood is derived by analyzing the structure of the $\eta_{k}$ in relation to the raw score and its derivatives $\partial_{\theta}^{k}\log f(\cdot;\theta)$. Importantly, note that we refer to results that use $\partial_{\theta}^{k}\log f(\cdot;\theta)$, or other derivatives such as the $\eta_{k}$, for $k\geq 1$ as Bhattacharyya-type bounds throughout the paper since the classical Bhattacharyya bound used $\frac{\partial^{k}_{\theta}f}{f}$ instead.
Our contributions are as follows.
  • We derive extrinsic geometric refinement terms using the second fundamental form in the non-asymptotic regime, providing insight into how curvature controls estimator error beyond linear approximations, and establish corrections to the classical CRB – see Theorem 2.
  • We interpret the higher-order variance lower bound sequence obtained by Bhattacharyya [8], and other authors following him, as successive $L^{2}$ projections, laying the foundation for geometric corrections making use of higher derivatives of the square root embedding and curvature information “beyond a certain order” (cf. Theorem 5). Thus, our version of the Bhattacharyya-type bound offers a clean geometric explanation for the differences between successive bounds.
  • By a careful analysis of the form of the jets $\eta_{k}$ expressed in terms of the raw score and its derivatives $\partial_{\theta}^{k}\log f(\cdot;\theta)$ via the Faà di Bruno formula and the exponential Bell polynomials, we identify orthogonal components arising from the corresponding expansions. It is established that $m$-th order Bhattacharyya-type bounds using log-likelihood derivatives are improved by those that use the jets $\{\eta_{k}\}_{k=1}^{m}$ when treating the aforementioned orthogonal components as extrinsic corrections; this result can be found in Proposition 8.
  • The proposed framework is exemplified by applying it to explicit models, demonstrating strict (and often fully analytical) improvement over both classical CRBs and algebraic higher-order bounds. Specifically, we endeavor to illuminate how curvature plays a role in quantifying estimator inefficiency.
To sum up/reiterate, this work unifies algebraic and geometric approaches to variance lower bounds and represents, to the best of our knowledge, the first use of extrinsic curvature to sharpen the CRB in the non-asymptotic scenario. It opens new avenues in the intersection of information geometry, Hilbert space theory, and statistical estimation.